OpenStudy (anonymous):

Solve the following equation for y 23x+7y=2

OpenStudy (anonymous):

You are given the following equation: \[23x + 7y = 2\] First we should try to move everything, but y-related things to one side. So, \(23x+7y = 2\) becomes \(7y = 2 - 23x\). Do you understand so far?

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

im not the best with math lol

OpenStudy (anonymous):

Okay, well keep note to this, when you move anything over the equal sign, if it is positive it becomes negative, if negative it becomes positive. So that is what I did above. Now that sort that out, we want \(y\) alone, so it is not with the constant \(7\). Do you have any idea on how to do that?

OpenStudy (anonymous):

oh okay and no I don't know how to do that

OpenStudy (anonymous):

You can divide, \(7y\) by \(7\) to get \(y\) by itself, however you have mirror that to the other side as well. (So you have to divide the other side by 7 as well.) \[\dfrac{7y}{7}=\dfrac{2 - 23x}{7} \implies y = \dfrac{2 - 23x}{7}\]

OpenStudy (anonymous):

oh okay now what ?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@Algorithmic

OpenStudy (anonymous):

Oh, sorry. That is your \(y\) nothing else to do, unless you want to graph it.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You are welcome.